OpenStudy (anonymous):

How do you convert 400 cm/sec to mi/hr?

OpenStudy (anonymous):

(400cm/sec)(1in/2.54cm)(1ft/12in)(1mi/5280ft) this is called dimensional analysis, do you see how the units cancel out?

OpenStudy (anonymous):

its easier if you write it like this \[(\frac{ 400cm }{\sec })(\frac{ 1 inch }{ 2.54cm })(\frac{ 1ft }{ 12inch })(\frac{ 1mi }{ 5280ft })(\frac{ 60\sec }{ 1\min })(\frac{ 60\min }{ 1hr })\]

OpenStudy (anonymous):

I see now thank you so much. I mean a lot!

OpenStudy (anonymous):

No problem! This was a big struggle for me in chem. I'm glad you see it now. If you have to take chemistry you can pretty much do magic if you understand DA